# Simple Estimation of Semiparametric Models with Measurement Errors

### v3

@@ 2.1 Special Case: Quadratic Expansion
where the second equality holds because because $\varepsilon_{i}$ and $\left(X_{i}^{\ast},S_{i}\right)$ are independent, and $\mathbb{E}\left[\varepsilon_{i}\right]=0$.
This expansion implies that $\mathbb{E}[g(X_{i},S_{i},\theta_{0})]=O\left(\sigma_{\varepsilon}^{2}\right)=O(\tau_{n}^{2})$. As a result, a naive estimator that ignores the EIV and uses $X_{i}$ in place of $X_{i}^{\ast}$ has EIV bias of order $\tau_{n}^{2}$. [^6] For example, consider a linear regression with a scalar mismeasured regressor. The bias of the naive OLS estimator of the slope parameter $\theta_{01}$ is $-\theta_{01}\frac{\tau_{n}^{2}}{1+\tau_{n}^{2}}=-\theta_{01}\tau_{n}^{2}+O\left(\tau_{n}^{4}\right)$. Bias of the naive estimator should be compared with its standard error, which is of order $n^{-1/2}$. Bias of the naive estimator is not negligible, unless the measurement error is rather small (theoretically, unless $\tau_{n}^{2}=o\left(n^{-1/2}\right)$). In particular, tests and confidence intervals based on the naive estimator are invalid and can provide highly misleading results. Moreover, if $\tau_{n}^{2}$ shrinks at a rate slower than $O\left(n^{-1/2}\right)$, the rate of convergence of the naive estimator is slower than $\sqrt{n}$.
@@
Here $O\left(\tau_{n}^{4}\right)=o\left(n^{-1/2}\right)$ because we assume that $\tau_{n}=o\left(n^{-1/6}\right)$. The idea behind this substitution is that the bias of order $O(\tau_{n}^{2})$ in $\mathbb{E}[g_{x}^{(2)}(X_{i},S_{i},\theta)]$ can be ignored because it is multiplied by $E\left[\varepsilon_{i}^{2}\right]=O\left(\tau_{n}^{2}\right)$. [^7] Such substitutions of $X^{\ast}$ with $X$ have been used in other contexts, e.g., Chesher and Schluter (2002). With the substitution, we can rearrange equation (4) and write it as
@@
where the first equality follows from equation (6) and the definition of $\gamma_{02}$, and the second equality follows from equation (1). Hence, the corrected moment conditions $\psi$ can be used to jointly estimate the true parameters $\theta_{0}$ and $\gamma_{02}$ by a GMM estimator. [^8] In the moment condition settings, having $o\left(n^{-1/2}\right)$ is equivalent to having 00 on the right-hand side of equation (8).
@@ 2.2 General Case: Expansion of order $K$
This equation is the general expansion analog of equation (4). The summation on the right hand side is the bias correction term, which we use to construct the MERM estimator.
@@
where $\gamma=(\gamma_{2},\dots,\gamma_{K})^{\prime}$ is a $K-1$ dimensional vector of parameters. Let $\gamma_{0}\equiv(\gamma_{02},\dots,\gamma_{0K})^{\prime}$ denote the vector of true parameters $\gamma_{0k}$, defined as
@@
$\varepsilon_{i}$ is independent from $(X_{i}^{\ast},S_{i})$ and $\mathbb{E}[\varepsilon_{i}]=0$.
Assumption 2.2 is the classical measurement error assumption. We relax this assumption later in Section 4.2. The following lemma establishes validity of the corrected moment conditions under Assumptions 2.1, 2.2, and some mild regularity conditions provided in Appendix A.
@@ 2.3 Measurement Error Robust Moments (MERM) estimator
$\beta\equiv(\theta^{\prime},\gamma^{\prime})^{\prime},\ \beta_{0}\equiv(\theta_{0}^{\prime},\gamma_{0}^{\prime})^{\prime},\text{ }\hat{\beta}\equiv(\hat{\theta}^{\prime},\hat{\gamma}^{\prime})^{\prime},$
@@
[…]
@@
Remark 3.
Researchers may be interested in average effects of the form $\lambda_{0}\equiv\mathbb{E}\left[\lambda\left(X_{i}^{\ast},S_{i},\theta_{0}\right)\right]$. In the NLR, one may be interested in the average partial effect $x$ (i.e., $\lambda_{0}\equiv\mathbb{E}\left[\nabla_{x}\rho\left(X_{i}^{\ast},S_{i},\theta_{0}\right)\right]$) or another covariate. The naive average partial effect estimator $\hat{\lambda}_{\text{Naive}}\equiv\frac{1}{n}\sum_{i=1}^{n}\lambda(X_{i},S_{i},\hat{\theta})$ suffers from the EIV bias, unless function $\lambda$ is linear in $X_{i}^{\ast}$. Instead, one can use estimates $\hat{\gamma}$ to construct the bias-corrected estimator $\hat{\lambda}_{\text{MERM}}\equiv\frac{1}{n}\sum_{i=1}^{n}\left\{\lambda(X_{i},S_{i},\hat{\theta})-\sum_{k=2}^{K}\hat{\gamma}_{k}\lambda_{x}^{(k)}(X_{i},S_{i},\hat{\theta})\right\}$.
Remark 4.
@@
It is important to note that $\gamma_{0k}\neq\mathbb{E}\left[\varepsilon_{i}^{k}\right]\left/k!\right.$ for $k\geq 4$, contrary to what equation (10) might suggest. For example, $\gamma_{04}=\left(\mathbb{E}\left[\varepsilon_{i}^{4}\right]-6\sigma_{\varepsilon}^{4}\right)\left/24\right.$ is negative for many distributions, including normal. For instance, in the example of Remark 4, $\gamma_{04}\approx-0.0026$. The reason that generally $\gamma_{0k}\neq\mathbb{E}\left[\varepsilon_{i}^{k}\right]\left/k!\right.$ is that the estimators of the correction terms themselves need a correction, which is accounted for by the form of $\gamma_{0k}$. Since there is a one-to-one relationship between $\gamma_{0}$ and the moments $\mathbb{E}\left[\varepsilon_{i}^{\ell}\right]$, parameter space $\Gamma$ for $\gamma_{0}$ can incorporate restrictions that the moments must satisfy (e.g., $\sigma_{\varepsilon}^{2}\geq 0$ and $\mathbb{E}\left[\varepsilon_{i}^{4}\right]\geq\sigma_{\varepsilon}^{4}$). Such restrictions can increase the efficiency of the estimator and the power of tests.
@@
Considering larger $K$ allows $\tau_{n}$converging to zero at a slower rate, which in finite samples corresponds to the asymptotics providing better approximations for larger magnitudes of measurement errors. On the other hand, taking a larger $K$ increases the dimension of the nuisance parameter $\gamma_{0}$ and thus typically increases the variance of $\hat{\theta}$.
Remark 8.
The formulas of the derivatives $g_{x}^{(k)}(\cdot)$ are typically easy to compute analytically or using symbolic algebra software. Alternatively, these derivatives can be computed using numerical differentiation. In either way, the corrected set of moments can be automatically produced for a generic moment function $g(\cdot)$ provided by the user.
@@ 2.4 Model Identification: Jacobian $\Psi$
for some moment function $u$. For example, consider the nonlinear regression model $\mathbb{E}\left[Y_{i}|X_{i}^{\ast}=x\right]=\rho\left(x,\theta_{0}\right)$, then $u\left(x,y,\theta\right)=\rho\left(x,\theta\right)-y$ [^9] For simplicity of the exposition, in the expectation in equation (15) we only condition on $X_{i}^{\ast}$. The discussion applies in a straightforward way to the settings with additional correctly measured variables $W_{i}$ in the conditioning set, i.e., the model $\mathbb{E}\left[u\left(X_{i}^{\ast},S_{i},\theta_{0}\right)|X_{i}^{\ast},W_{i}\right]=0$. For example, in the nonlinear regression example with additional covariates $W_{i}$ we have $\mathbb{E}\left[Y_{i}|X_{i}^{\ast}=x,W_{i}=w\right]=\rho\left(x,w,\theta_{0}\right)$, so $u\left(x,y,w,\theta\right)=\rho\left(x,w,\theta\right)-y$.
@@
We attempt to address this concern and make the above intuitions more precise in two ways. First, in the following subsection we consider a specific (but frequently employed) kind of an instrument: a second measurement (possibly nonclassical). The specific form of the excluded variable allows us to provide more transparent identification conditions. Second, in Evdokimov and Zeleneev (2022) we study nonparametric regression model with EIV using the $\tau_{n}\rightarrow 0$ approximation. We show that the model is identified using an instrument (even a discrete one), and motivate the MERM approach from a nonparametric perspective. Finally, since MERM estimator is a standard GMM estimator, one can test the strength of identification of the model parameters, or conduct identification-robust inference using the standard methods (e.g., StockWright2000Ecta,Kleibergen2005Ecta,GuggenbergerSmith2005ET,GuggenbergerRamalhoSmith2012JoE,AndrewsMikusheva2016Ecta-ConditionalFunctionalNuisance,AndrewsI-2016-Ecta-CLC,AndrewsGuggenberger2019QE).
@@
It turns out that in these settings there is a simple sufficient condition for Assumption 2.3(i) to hold. Appendix D demonstrates that $\Psi^{\ast}$ will have full rank if
@@
2.5 Assessing model validity
The standard $J$-test of overidentifying restrictions remains valid in the MERM settings, and can be used to check the model specification. The $J$-test jointly tests the following hypotheses: (i) $K$ is sufficiently large to correct the EIV bias; (ii) assumptions on the EIV are valid; and (iii) the original moment conditions $g$ are correctly specified so equation (1) holds, i.e., that the original economic model is correctly specified aside from the presence of the EIV in $X_{i}$. Issue (i) can be address by taking a larger $K$. In Section 4 we extend the framework to obtain corrected moments that are valid under weaker assumptions on the EIV, which can help addressing issue (ii).
It is possible to obtain an explicit bound on the higher-order EIV bias of the MERM estimator. Suppose the researcher is interested in estimating a linear combination of parameters $v^{\prime}\beta_{0}$ for some non-zero $v\in\mathbb{R}^{\dim(\beta)}$, e.g., the first component of $\theta_{0}$. Since $\hat{\beta}$ is a standard GMM estimator,
$\text{Bias}(v^{\prime}\hat{\beta})\approx-v^{\prime}(\Psi^{\prime}\Xi\Psi)^{-1}\Psi^{\prime}\Xi\;\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})\right].$ (18)
Here $\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})\right]\approx\gamma_{0\overline{K}}\mathbb{E}[g_{x}^{(\overline{K})}(X_{i},S_{i},\theta_{0})]$ with $\overline{K}=K+2$ if $K$ is even and $\varepsilon_{i}$ is symmetric, and $\overline{K}=K+1$ otherwise. Thus,
$\text{Bias}(v^{\prime}\hat{\beta})\approx\gamma_{0\overline{K}}b_{v},$
where $b_{v}$ can be easily estimated. The value of $\gamma_{0\overline{K}}$ is given by the recursive relationship (12) and determined by the unknown $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ (and by the lower order moments of $\varepsilon_{i}$, which can be recovered from $\hat{\gamma}$ with a sufficient precision). While $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ and, hence, $\gamma_{0\overline{K}}$ are generally unknown, one simple way to evaluate the magnitude of $\text{Bias}(v^{\prime}\hat{\beta})$ is to calculate $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ using the estimated $\sigma_{\varepsilon}^{2}$ and assuming normality of $\varepsilon$. Alternatively, one can also bound $\gamma_{0\overline{K}}$ and, subsequently, $\text{Bias}(v^{\prime}\hat{\beta})$ by imposing some (a priori) bounds on $\mathbb{E}[(\varepsilon_{i}/\sigma_{\varepsilon})^{\overline{K}}]$.
The bounds on Bias$(v^{\prime}\hat{\beta})$ can also be used to assess the appropriateness of Assumption 2.1 and the reliability of the inference based on the normal approximation in equation (14). If the (worst case) bias of $v^{\prime}\hat{\beta}$ is substantial, relative to its standard error, the researcher may want to account for this and consider choosing a higher $K$. Appendix E provides a further discussion of this issue and additional details on the calculation of Bias$(v^{\prime}\hat{\beta})$.
@@ 3.1 Comparison with a Semi-Nonparametric Estimation Approach
We compare MERM estimator with the state-of-the-art semiparametric estimator of Schennach (2007, henceforth S07) for nonlinear regression models. The Monte Carlo designs are taken from S07, and include a polynomial, rational fraction, and Probit nonlinear regression models. Identification of the model is ensured by the availability of an instrument.
@@
$(Z_{i},V_{i},\varepsilon_{i})^{\prime}\sim N\left((0,0,0)^{\prime},\text{Diag}(1,1/4,1/4)\right)$, $\pi_{1}=1$, and $n=1000$. The conditional expectation function $\rho$, the true value of the parameter of interest $\theta_{0}$, and the conditional distribution of the regression error $U_{i}$ are design-specific and reported in Tables 1-3 below. In all designs, $\tau=\sigma_{\varepsilon}/\sigma_{X}^{*}\approx 0.45$, so the measurement error is “fairly large” (Schennach, 2007).
@@
$K=4$0.01 -0.01 0.23 0.42 0.23 0.42 0.48
  • The DGP is as in (19) with $\rho(x,\theta)=\frac{1}{2}(1+\text{erf}(\theta_{1}+\theta_{2}x))$, $\theta_{0}=(-1,2)^{\prime}$, and $U_{i}=1-\rho(X_{i}^{*},\theta_{0})$ with probability $\rho(X_{i}^{*},\theta_{0})$ and $-\rho(X_{i}^{*},\theta_{0})$ otherwise.

### v4

@@ 2.1 Special Case: Quadratic Expansion
where the second equality holds because $\varepsilon_{i}$ and $\left(X_{i}^{\ast},S_{i}\right)$ are independent, and $\mathbb{E}\left[\varepsilon_{i}\right]=0$}.
Evaluating the expansion above at $\theta=\theta_{0}$ gives $\mathbb{E}[g(X_{i},S_{i},\theta_{0})]=O\left(\sigma_{\varepsilon}^{2}\right)=O(\tau_{n}^{2})$, because $\mathbb{E}[g(X_{i}^{*},S_{i},\theta_{0})]=0$. As a result, a naive estimator that ignores the EIV and uses $X_{i}$ in place of $X_{i}^{\ast}$ has EIV bias of order $\tau_{n}^{2}$. [^6] For example, consider a linear regression with a scalar mismeasured regressor. The bias of the naive OLS estimator of the slope parameter $\theta_{01}$ is $\theta_{01}\frac{\tau_{n}^{2}}{1+\tau_{n}^{2}}=-\theta_{01}\tau_{n}^{2}+O\left(\tau_{n}^{4}\right)$. The bias of the naive estimator should be compared with its standard error, which is of order $n^{-1/2}$. Thus, the bias of the naive estimator is not negligible, unless the measurement error is rather small (theoretically, unless $\tau_{n}^{2}=o\left(n^{-1/2}\right)$). In particular, tests and confidence intervals based on the naive estimator are invalid and can provide highly misleading results. Moreover, if $\tau_{n}^{2}$ shrinks at a rate slower than $O\left(n^{-1/2}\right)$, the rate of convergence of the naive estimator is slower than $\sqrt{n}$.
@@
Here $O\left(\tau_{n}^{4}\right)=o\left(n^{-1/2}\right)$ because we assume that $\tau_{n}=o\left(n^{-1/6}\right)$. The idea behind this substitution is that the bias of order $O(\tau_{n}^{2})$ in $\mathbb{E}[g_{x}^{(2)}(X_{i},S_{i},\theta)]$ can be ignored because it is multiplied by $E\left[\varepsilon_{i}^{2}\right]=O\left(\tau_{n}^{2}\right)$. [^7] Such substitutions of $X^{\ast}$ with $X$ have been used in other contexts, e.g., ChesherSchluter2002ReStud. With the substitution, we can rearrange equation (4) and write it as
@@
where the first equality follows from equation (6) and the definition of $\gamma_{02}$, and the second equality follows from equation (1). Hence, the corrected moment conditions $\psi$ can be used to jointly estimate the true parameters $\theta_{0}$ and $\gamma_{02}$ by a GMM estimator. [^8] In the moment condition settings, having $o\left(n^{-1/2}\right)$ is equivalent to having 0 on the right-hand side of equation (8).
@@ 2.2 General Case: Expansion of order $K$
[…]
@@
where $\gamma=(\gamma_{2},\dots,\gamma_{K})^{\prime}$ is a K1K-1 dimensional vector of parameters. Let $\gamma_{0}\equiv(\gamma_{02},\dots,\gamma_{0K})^{\prime}$ denote the vector of true parameters $\gamma_{0k}$, defined as
@@
(Classical Measurement Error) $\varepsilon_{i}$ is independent from $(X_{i}^{\ast},S_{i})$ and $\mathbb{E}[\varepsilon_{i}]=0$.
The following lemma establishes validity of the corrected moment conditions under Assumptions 2.1, 2.2, and some mild regularity conditions provided in Appendix A.
@@
Remark 2.
Construction of the corrected moment conditions $\psi$ requires the original moment function $g(x,s,\theta)$ to have a sufficient number of derivatives with respect to $x$. Thus, the proposed correction method does not apply to settings with non-differentiable moment functions, for example, those arising in the instrumental variable quantile regression (IVQR).
@@ 2.3 Measurement Error Robust Moments (MERM) estimator
[…]
@@
$n^{1/2}\Sigma^{-1/2}(\hat{\beta}-\beta_{0})\overset{d}{\rightarrow}N(0,I_{\dim(\beta)}),\text{ where}$ (14)
@@
The standard $J$-test of overidentifying restrictions remains valid in the MERM settings, and can be used to check the model specification. The $J$-test jointly tests the following hypotheses: (i) $K$ is sufficiently large to correct the EIV bias; (ii) assumptions on the EIV are valid; and (iii) the original moment conditions $g$ are correctly specified so equation (1) holds, i.e., that the original economic model is correctly specified aside from the presence of the EIV in $X_{i}$. Thus, if the $J$-test rejects the validity of the corrected moment conditions $\psi$, the researcher might want to (i) consider taking a larger $K$; (ii) employ a different correction method; or (iii) consider an alternative specification of the original moments $g$.
Remark 5.
Considering larger $K$ allows for $\tau_{n}\$converging to zero at a slower rate, which in finite samples corresponds to the asymptotics providing better approximations for larger magnitudes of measurement errors. On the other hand, taking a larger $K$ increases the dimension of the nuisance parameter $\gamma_{0}$ and thus typically increases the variance of $\hat{\theta}$. We consider this issue in more detail and provide a data-driven method for choosing $K$ in Section 4.1.
Remark 6.
The MERM framework can be extended to the case of non-classical measurement errors; see EvdokimovZeleneev2022WP-NPID for details and a fully nonparametric analysis. In this paper, we focus on the classical measurement errors, developing a practical bias correction approach, which can be easily implemented in a wide range of economic applications. Even when Assumption 2.2 is violated, the deviations from it are often limited in magnitude, so the corrections based on the 2.2 assumption remove most of the EIV bias. Thus, in practice, using the estimator designed for classical measurement errors is typically preferable to ignoring mismeasurement altogether.
Remark 7.
It is important to note that $\gamma_{0k}\neq\mathbb{E}\left[\varepsilon_{i}^{k}\right]\left/k!\right$. for $k\geq 4$, contrary to what equation (10) might suggest. For example, $\gamma_{04}=\left(\mathbb{E}\left[\varepsilon_{i}^{4}\right]-6\sigma_{\varepsilon}^{4}\right)\left/24\right$. is negative for many distributions, including normal. The reason that generally $\gamma_{0k}\neq\mathbb{E}\left[\varepsilon_{i}^{k}\right]\left/k!\right$. is that the estimators of the correction terms themselves need a correction, which is accounted for by the form of $\gamma_{0k}$. Since there is a one-to-one relationship between $\gamma_{0}$ and the moments $\mathbb{E}\left[\varepsilon_{i}^{\ell}\right]$, parameter space $\Gamma$ for $\gamma_{0}$ can incorporate restrictions that the moments must satisfy (e.g., $\sigma_{\varepsilon}^{2}\geq 0$ and $\mathbb{E}\left[\varepsilon_{i}^{4}\right]\geq\sigma_{\varepsilon}^{4}$). Such restrictions can increase the efficiency of the estimator and the power of tests.
@@
The formulas of the derivatives $g_{x}^{(k)}(\cdot)$ are typically easy to compute analytically or using symbolic algebra software. Alternatively, these derivatives can be computed using numerical differentiation. Thus, the corrected set of moments can be automatically produced for a generic moment function $g(\cdot)$ provided by the user.
@@ 2.4 Model Identification: Jacobian $\Psi$
for some moment function $u$. For example, consider the nonlinear regression model $\mathbb{E}\left[Y_{i}|X_{i}^{\ast}=x\right]=\rho\left(x,\theta_{0}\right)$, then $u\left(x,y,\theta\right)=\rho\left(x,\theta\right)-y$. [^10] For simplicity of the exposition, in the expectation in equation (15) we only condition on $X_{i}^{\ast}$. The discussion applies in a straightforward way to the settings with additional correctly measured variables $W_{i}$ in the conditioning set, i.e., the model $\mathbb{E}\left[u\left(X_{i}^{\ast},S_{i},\theta_{0}\right)|X_{i}^{\ast},W_{i}\right]=0$. For example, in the nonlinear regression example with additional covariates $W_{i}$ we have $\mathbb{E}\left[Y_{i}|X_{i}^{\ast}=x,W_{i}=w\right]=\rho\left(x,w,\theta_{0}\right)$, so $u\left(x,y,w,\theta\right)=\rho\left(x,w,\theta\right)-y$.
@@
We attempt to address this concern and make the above intuitions more precise in two ways. First, in the following subsection we consider a specific (but frequently employed) kind of an instrument: a second measurement (possibly non-classical). The specific form of the excluded variable allows us to provide more transparent identification conditions. Second, in EvdokimovZeleneev2022WP-NPID we study nonparametric regression model with EIV using the $\tau_{n}\rightarrow 0$ approximation. We show that the model is identified using an instrument (even a discrete one), and motivate the MERM approach from a nonparametric perspective. Finally, since MERM estimator is a standard GMM estimator, one can test the strength of identification of the model parameters, or conduct identification-robust inference using the standard methods (e.g., StockWright2000Ecta,Kleibergen2005Ecta,GuggenbergerSmith2005ET,GuggenbergerRamalhoSmith2012JoE,AndrewsMikusheva2016Ecta-ConditionalFunctionalNuisance,AndrewsI-2016-Ecta-CLC,AndrewsGuggenberger2019QE).
@@
It turns out that in these settings there is a simple sufficient condition for Assumption 2.3(i) to hold. Appendix E demonstrates that $\Psi^{\ast}$ will have full rank if
@@ 3.1 Comparison with a Semi-Nonparametric Estimation Approach
We compare MERM estimator with the state-of-the-art semiparametric estimator of Schennach2007Ecta for nonlinear regression models. The Monte Carlo designs are taken from S07, and include a polynomial, rational fraction, and Probit nonlinear regression models. Identification of the model is ensured by the availability of an instrument.
@@
$(Z_{i},V_{i},\varepsilon_{i})^{\prime}\sim N\left((0,0,0)^{\prime},\text{Diag}(1,1/4,1/4)\right)$, $\pi_{1}=1$, and $n=1000$. The conditional expectation function $\rho$, the true value of the parameter of interest $\theta_{0}$, and the conditional distribution of the regression error $U_{i}$ are design-specific and reported in Tables 1-3 below. In all designs, $\tau=\sigma_{\varepsilon}/\sigma_{X}^{*}\approx 0.45$, so the measurement error is “fairly large” (Schennach2007Ecta).
@@
$K=4$ -0.01 -0.01 0.23 0.42 0.23 0.42 0.48
  • The DGP is as in (18) with $\rho(x,\theta)=\frac{1}{2}(1+\text{erf}(\theta_{1}+\theta_{2}x))$, $\theta_{0}=(-1,2)^{\prime}$, and $U_{i}=1-\rho(X_{i}^{*},\theta_{0})$ with probability $\rho(X_{i}^{*},\theta_{0})$ and $\rho(X_{i}^{*},\theta_{0})$ otherwise.